OpenStudy (cas_fangirl_14):

WILL MEDAL 6*8+2=? DISCLAMER: This is not a medal party of any kind. I just want reward people for giving good answers.

OpenStudy (wolfman77):

Use PEMDAS. 6*8+2 6 * 8 = 48 + 2 = 50. :) Hope this Helps!

OpenStudy (cas_fangirl_14):

Thanks

OpenStudy (koolkid98):

it 50 ^

OpenStudy (wolfman77):

Mhm. No Problem!

OpenStudy (koolkid98):

wolfman is right